# Forecasting high-impact research topics via machine learning on evolving knowledge graphs

### 13 Feb 2024 v1

@@ Introduction
An alternative and complementary approach is to build scientific semantic knowledge graphs. Here, the nodes represent scientific concepts and the edges are formed when two concepts are researched together in a scientific paper [2]. While this approach extracts only small amounts of information from each paper, surprisingly non-trivial conclusions can still be drawn if the underlying dataset of papers is large. An early example of this is a work in biochemistry [9]. The authors use their semantic network, where nodes represent biomolecules, to find new potentially more efficient exploration strategies for the bio-chemistry community on a global scale. In these semantic networks, an edge between two concepts indicates that researchers have jointly investigated these research concepts. The edges are drawn from papers, thus they are created at a specific time when the paper was published. In this way, one creates an evolving semantic network that captures what researchers have investigated in the past. With such an evolving network, one can ask how the network might evolve in the future. In the scientific context, this question can be reformulated into what scientists will research in the future. For example, if two nodes do not share an edge, one can ask whether they will share an edge in the next three years – or, alternatively, whether scientists will investigate these two concepts jointly within three years. This question, denoted as link-prediction problem in network theory [10], has been successfully demonstrated with high prediction quality for semantic networks in quantum theory [11] and artificial intelligence [4].
@@
To build our concept list, we started with 2,444,442 papers from four publicly available preprint servers: arXiv, bioRxiv, medRxiv, and ChemRxiv. The data cutoff is February 2023. From these, we extracted titles and abstracts of the papers. To single out concept candidates from this extensive collection, we applied the Rapid Automatic Keyword Extraction (RAKE) algorithm based on statistical text analysis to automatically detect important keywords [41]. Concepts with two words, like phase transition, were retained if they appear in at least 9 papers, while longer concepts, such as single molecule localization microscopy, needed to appear in at least 6 papers. In this way, we can increase the ratio of high-quality concepts. To refine our list further, we developed a suite of natural language processing tools. Finally, we got a list which contains over 368,000 concepts. While we focus here on concepts specific to the sub-field of optics and quantum physics (representing roughly 10% of the entire concepts), our method can immediately be translated to any other domain. This refined domain-specific concept list serves as the vertices for our knowledge graph.
@@
Creating an evolving, citation-augmented knowledge graph – Now that we have the vertices, we can create edges that contain information from the scientific literature. To have citation information, we use the works from OpenAlex [42], an open-source database containing detailed information on more than 90 million scientific publications. Edges are drawn when two concepts co-occur in the title or abstract of a scientific paper. If a paper connects two vertices, the weight of the newly formed edge is the paper’s annual citation numbers from 2012 to 2023 together with the total citation number since its publication. If more than one paper creates an edge, then the edge contains the sum of the annual citations (as well as the sum of the total citations) gained by all papers. As research papers appear over time, and their citations are created in time, we effectively build an evolving, citation-augmented knowledge graph that evolves in time (see Fig. 1).
@@
The final constructed knowledge graph has 37,960 vertices with more than 26 million edges (built from 190 million concept pairs, containing multi-edges when multiple papers create the same edge) from the OpenAlex database, with a data cutoff at April 2023. In Fig. 2, we see that the distribution of the 3-year citation increase for previously uncited concept pairs exhibits a heavy tail. This suggests that some concept pairs are cited significantly more than would be expected from an exponential decrease, potentially due to the influence of concept hubs [4]. In Fig. 3, we show the fastest growing (in terms of citation) concepts and concept pairs since 2012, where we can recognize many highly influential topics in quantum physics and optics research.
Results
@@
Forecasting impact of newly created concept connections – With an evolving knowledge graph from the past, we can formulate the prediction of impact for new concept pairs as a supervised learning task, as illustrated in Fig. 4. For a vertex pair that has not had any connection in the year 2016, we predict whether 3 years later this vertex pair accumulated more than a certain number of citations. Using the historical knowledge graph, we possess an ideal supervision signal for our binary classification task. During the training phase, we selected pairs of vertices that were not connected and calculated 141 features for each pair. These features include 41 network features, divided into 20 node features (such as the number of neighbors and PageRank [43] over the past three years) and 21 edge features (including cosine, geometric, and Simpson similarities [44]). Additionally, we incorporated 100 impact features: 58 of these are node citation features, covering total citations and the count of papers mentioning the concept within the last three years. The other 42 features are about vertex pairs and include measures such as the citation ratio between them. The network feature is inspired by the winner of the Science4Cast competition [45, 11], and the citation features are developed empirically and could potentially be improved by careful feature importance analysis. Our neural network is a fully connected feed-forward network with four hidden layers of 600 neurons each. The exploration of more advanced architectures might improve the prediction qualities further. The neural network has to predict whether the unconnected vertex pair in 2019 will have more than citations ( is impact range).
@@
We perform the training from to . We quantify the quality with the area under curve (AUC) of the receiver operating characteristic curve (ROC) [46]. The AUC gives a measure of classification quality and stands for the probability that a randomly chosen true example is ranked higher than a randomly chosen false example. A random classifier has . We measure the AUC for a test set (which contains unconnected pairs not in the training set) for a prediction from 2016 to 2019, and for an evaluation dataset, with 10 million random data from 2019 to 2022 (while keeping the training data of the neural network from the year 2016 to 2019). The evaluation dataset shows how well the neural network performs on future, never-seen datasets. This is motivated by our goal that ultimately we want to train a neural network with all available data (let’s say, until the end of 2023) and predict what happens until the future in 2026. In Fig. 5(a), we find that the AUC scores for both the test set and the evaluation set are beyond 0.8, in most of the cases beyond 0.9, for different values of the impact range . We can conclude that the neural network can forecast a high impact of previously never-investigated concept connections to a high degree. In Fig. 5(b), we sort the concept pairs of the evaluation dataset with the neural network (), and plot their true citation counts. We further divide the 10 million evaluation dataset into 20 equal parts and plot their average citation count (represented by green bars) for each 5% segment. This clearly demonstrates very good predictions at the individual concept pair level.
@@
Forecasting genuine impact beyond link prediction – Next, we perform an even more challenging, genuine impact prediction task that goes beyond link prediction (i.e., predicting which concept pairs will be investigated in the future by a scientific paper). Concretely, in this second task training data is conditioned on unconnected vertex pairs from 2016 which are actually connected in 2019. The neural network only gets information from 2016 and has to predict whether the newly generated concept pair will be highly impactful or not. For that, our classification task asks whether the newly generated edge will receive between 0-5 citations or above 100 (Fig. 5(c)). We see that the AUC is beyond 0.7 (for the test set) and beyond 0.67 for the evaluation set, clearly indicating that the neural network can predict impact properties that go beyond the simpler link-prediction task.
Extracting highly impactful cliques of concepts – In what follows, we show one way how our results can be applied to inspire new directions that potentially human researchers would not have thought about. So far, we have limited our predictions to pairs of concepts. However, incorporating a larger number of concepts could more directly indicate a research direction. Directly predicting high-impact concept triples or quadruples is either highly computationally expensive or necessitates entirely different data structures, such as hypergraphs. Alternatively, we can approximate multi-concept combinations in our knowledge graph by searching high-impact cliques. Cliques are fully connected subgraphs with N vertices. If , we have concept pairs, but for higher , we get a larger number of concepts.
To get the cliques, we train a neural network to predict high-impact (for example, ) with all available data and predict what currently unconnected pairs will be highly impactful until 2026. We then apply the neural network to all unconnected vertex pairs (694 million pairs) and sort the result from highly likely to be of high impact, to least likely. With this sorted edge list, we create a subgraph, starting from the highest predicted edge (cancer cell with renewable energy) and add edges one by one. In Fig. 6, we collect the first cliques of size ranging from 2 to 8. We eagerly await the year 2026 to see whether a paper investigating the concepts Hilbert space, QR code, transformer model and bibliometric analysis will indeed be highly impactful. In a real application scenario, one could personalize the list of concepts to fit the research interest of individual scientists.
Outlook
@@
We show how to forecast the impact of future research topics. Although we view this as a significant step towards developing truly useful AI-driven assistants, achieving this goal requires numerous further advancements. Firstly, developing methods to extract more complex information from each paper will be crucial, for instance by employing hyper-graph structures that carry more information from each paper [47, 48], which has already been demonstrated to lead to exciting results in other domains[49, 2, 50]. This might also allow for the forecast of new concepts [51, 52] and their impact. Secondly, it will be interesting to approximate impact with metrics that go beyond citations – which is a crucial topic in computational sociology and the study of the science of science [1, 2]. Additionally, introducing metrics of surprise, as discussed in [53, 54], could serve as a complementary metric to citation prediction for ranking suggestions. Finally, while the suggestion of impactful new ideas might be a crucial component of future AI assistants, it will be crucial to study its relation to the scientific interest of working researchers.

### 03 Mar 2024 v2

@@ Introduction
An alternative and complementary approach is to build scientific semantic knowledge graphs. Here, the nodes represent scientific concepts and the edges are formed when two concepts are researched together in a scientific paper [2]. While this approach extracts only small amounts of information from each paper, surprisingly non-trivial conclusions can be drawn if the underlying dataset of papers is large. An early example of this is a work in biochemistry [9]. The authors use their semantic network, where nodes represent biomolecules, to find new potentially more efficient exploration strategies for the bio-chemistry community on a global scale. In these semantic networks, an edge between two concepts indicates that researchers have jointly investigated these research concepts. The edges are drawn from papers, thus they are created at a specific time when the paper was published. In this way, one creates an evolving semantic network that captures what researchers have investigated in the past. With such an evolving network, one can ask how the network might evolve in the future. In the scientific context, this question can be reformulated into what scientists will research in the future. For example, if two nodes do not share an edge, one can ask whether they will share an edge in the next three years – or, alternatively, whether scientists will investigate these two concepts jointly within three years. This question, denoted as link-prediction problem in network theory [10], has been successfully demonstrated with high prediction quality for semantic networks in quantum theory [11] and artificial intelligence [4].
@@ Results
To build our concept list, we started with 2,444,442 papers from four publicly available preprint servers: arXiv, bioRxiv, medRxiv, and chemRxiv. The data cutoff is February 2023. From these, we extracted titles and abstracts of the papers. To single out concept candidates from this extensive collection, we applied the Rapid Automatic Keyword Extraction (RAKE) algorithm based on statistical text analysis to automatically detect important keywords [36]. Concepts with two words, like phase transition, were retained if they appear in at least 9 papers, while longer concepts, such as single molecule localization microscopy, needed to appear in at least 6 papers. In this way, we can increase the fraction of high-quality concepts. To refine our list further, we developed a suite of natural language processing tools. Finally, we got a list which contains over 368,000 concepts. We focus here on concepts specific to the sub-field of optics and quantum physics (representing roughly 10% of the entire concepts), but our method can immediately be translated to any other domain. This refined domain-specific concept list serves as the vertices for our knowledge graph.
@@
Creating an evolving, citation-augmented knowledge graph – Now that we have the vertices, we can create edges that contain information from the scientific literature. We get citation information from papers in OpenAlex [37], an open-source database containing detailed information on more than 90 million publications. Edges are drawn when two concepts co-occur in the title or abstract of a scientific paper. If a paper connects two vertices, the weight of the newly formed edge is the paper’s annual citation numbers from 2012 to 2023 together with the total citation number since its publication. If more than one paper creates an edge, then the edge contains the sum of the annual citations (as well as the sum of the total citations) gained by all papers. As research papers appear over time, and their citations are created in time, we effectively build an evolving, citation-augmented knowledge graph that evolves in time (see Fig. 1).
@@
The final constructed knowledge graph has 37,960 vertices with more than 26 million edges (built from 190 million concept pairs, containing multi-edges when multiple papers create the same edge) from the OpenAlex database, with a data cutoff at April 2023. In Fig. 2, we see that the distribution of the 3-year citation increase for previously uncited concept pairs exhibits a heavy tail. This suggests that some concept pairs are cited significantly more than would be expected from an exponential decrease. In Fig. 3, we show the fastest growing (in terms of citation) concepts and concept pairs since 2012, where we can recognize many highly influential topics in quantum physics and optics research.
@@
Forecasting impact of newly created concept connections – With an evolving knowledge graph from the past, we can formulate the prediction of impact for new concept pairs as a supervised learning task, as illustrated in Fig. 4. For a vertex pair that has not had any connection in the year 2016, we predict whether 3 years later this vertex pair accumulated more than a certain number of citations. Using the historical knowledge graph, we possess an ideal supervision signal for our binary classification task. During the training phase, we selected pairs of vertices that were not connected and calculated 141 features for each pair. These features include 41 network features, divided into 20 node features (such as the number of neighbors and PageRank [43] over the past three years) and 21 edge features (including cosine, geometric, and Simpson similarities [44]). Additionally, we incorporated 100 impact features: 58 of these are node citation features, covering total citations and the count of papers mentioning the concept within the last three years. The other 42 features are about vertex pairs and include measures such as the citation ratio between them. The network features are inspired by the winner of the Science4Cast competition [45, 11], and the citation features are developed empirically and could potentially be improved by careful feature importance analysis. Our neural network is a fully connected feed-forward network with four hidden layers of 600 neurons each. The exploration of more advanced architectures might improve the prediction qualities further. The neural network has to predict whether the unconnected vertex pair in 2019 will have at least citations ( is impact range).
@@
We perform the training from to . We quantify the quality with the area under curve (AUC) of the receiver operating characteristic curve (ROC) [46]. The AUC gives a measure of classification quality and stands for the probability that a randomly chosen true example is ranked higher than a randomly chosen false example. A random classifier has . We measure the AUC for a test set (which contains unconnected pairs not in the training set) for a prediction from 2016 to 2019, and for an evaluation dataset, with 10 million random data from 2019 to 2022 (while keeping the training data of the neural network from the year 2016 to 2019). The evaluation dataset shows how well the neural network performs on future, never-seen datasets. This is motivated by our goal that ultimately we want to train a neural network with all available data (let’s say, until January 2023) and predict what happens until the future in 2026. In Fig. 5(a), we find that the AUC scores for both the test set and the evaluation set are beyond 0.8, in most of the cases beyond 0.9, for different values of the impact range . We can conclude that the neural network can forecast a high impact of previously never-investigated concept connections to a high degree. In Fig. 5(b), we sort the concept pairs of the evaluation dataset with the neural network (), and plot their true citation counts. We further divide the 10 million evaluation dataset into 20 equal parts and plot their average citation count (represented by green bars) for each 5% segment. This clearly demonstrates very good predictions at the individual concept pair level. As seen in Fig. 5(c), the highest predicted concept pairs indeed get more than 5 orders of magnitude more citations than the average citation of all 10 pairs.
@@
Forecasting genuine impact beyond link prediction – Next, we perform an even more challenging, genuine impact prediction task that goes beyond link prediction (i.e., predicting which concept pairs will be investigated in the future by a scientific paper). Concretely, in this task training data is conditioned on unconnected vertex pairs from 2016 which are actually connected in 2019. The neural network only gets citation information from 2016 and has to predict whether the newly generated concept pair will be highly impactful or not. For that, our classification task asks whether the newly generated edge will receive between 0-5 citations or above 100 (Fig. 5(d)). We see that the AUC is beyond 0.7 (for the test set) and beyond 0.67 for the evaluation set, clearly indicating that the neural network can predict impact properties that go beyond the simpler link-prediction task.
Highly predicted impact pair and potential applications – We can now investigate the largest predicted pairs of concepts, by taking all unconnected vertex pairs (694 million pairs) until 2023, and let the neural network (trained with all data until 2023) sort them by impact. We find that the highest predicted pair is renewable energy and cancer cell. This prediction is a very high-risk bet. For more practical, personalized suggestions, one can restrict the unconnected concept pairs to those related to specific scientists or research groups, aiming for high-impact collaboration suggestions. By examining the published works of scientists to identify their research interests, it becomes possible to identify concept pairs where one aligns with one scientist’s specialty and the other with another scientist’s. Thereby, one can suggest potential collaborations of significant impact. As an example, by constraining the personalized research interests of scientists in experimental quantum optics and one researcher in biophysics, the highest predicted impact concepts pairs are ‘microfluidic channel’ with ‘Kerr resonator’, ‘SARS CoV’ with ‘quantum enhanced sensitivity’ or ‘electron microscopy’ with ‘quantum vacuum field’. These suggestions can be further refined based on their similarity (e.g., represented by the cosine similarity) or the prominence of the concepts (indicated by the node degree), as we show in Fig. 6. Here, we plot 100,000 concept pairs that have not been studied together until January 2023 and use a neural network trained on 2019 dataset to predict their impact. The points are plotted based on various properties, such as the similarity between concepts, their prominence within the network, their growth rate in the network (reflected in newly acquired neighbors), and how often the concepts have been cited previously. Plotting in this way allows us to identify rare outliers – concept pairs with high predicted impact that have unique properties, such as the bright yellow spots highlighted in the insets of Fig. 6. These methods help us narrow down the enormously large number of possibilities into a small number of personalized and targeted suggestions, which could inspire new ideas.
Discussion
@@
We show how to forecast the impact of future research topics. Although we view this as a significant step towards developing truly useful AI-driven assistants, achieving this goal requires numerous further advancements. Firstly, developing methods to extract more complex information from each paper will be crucial, for instance by employing hyper-graph structures that carry more information from each paper [47, 48], which has already been demonstrated to lead to exciting results in other domains[49, 2, 50]. This might also allow for the forecast of new concepts [51, 52] and their impact. Secondly, it will be interesting to approximate impact with metrics that go beyond citations – which is a crucial topic in computational sociology and the study of the science of science [1, 2]. Additionally, introducing metrics of surprise, as discussed in [53, 54], could serve as a complementary metric to citation prediction for ranking suggestions. Finally, while the suggestion of impactful new ideas might be a key component of future AI assistants, it will be crucial to study its relation to the scientific interest of working researchers.
Methods
Datasets for knowledge graph construction – To compile a list of scientific concepts in natural science, we used metadata from four major preprint servers: arXiv, bioRxiv, medRxiv, and chemRxiv. The arXiv dataset can be directly downloaded from Kaggle, while metadata from bioRxiv, medRxiv, and chemRxiv are accessible through their APIs. The full methodology and code are available on the GitHub: Impact4Cast. Our comprehensive dataset encompasses approximately 2.44 million papers, including 78,084 from arXiv’s physics.optics and quant-ph categories, which were specifically utilized for identifying domain concepts.
For edge generation, we used the OpenAlex database snapshot, available for download in OpenAlex bucket stored in Amazon S3. More details can be found at the OpenAlex documentation site. The complete dataset occupies around 330 GB, expanding to approximately 1.6 TB when decompressed. Our interest was specifically in journal papers that include publication time, title, abstract, and citation information. By focusing on these criteria, we managed to reduce the dataset to a more manageable gzip-compressed size of 68 GB, comprising around 92 million scientific papers.
Details on concept and edge generation – From the preprint dataset of 2.44 million papers, we analyzed each article’s title and abstract using the RAKE algorithm, enhanced with additional stopwords, to extract potential concept candidates. These candidates were stored for subsequent analysis. We filtered out concepts to retain only those with two words that appeared in nine or more articles, and those with three or more words that appeared in six or more articles. This step significantly reduced the noise from the RAKE-generated concepts, yielding a refined list of 726,439 relevant concepts. To further improve the quality of these concepts, we developed an automatic tool to identify and remove common, domain-independent errors often made by RAKE. This process, which included eliminating non-conceptual phrases, verbs, ordinal numbers, conjunctions, and adverbials, resulted in a full list of 388,825 concepts.
We then specifically focused on articles within the physics.optics and quant-ph categories from arXiv to extract domain-specific concepts. By iterating our entire list of concepts to these domain-specific articles, we identified 87,741 relevant concepts. Employing our specially designed automated filtering tool for further refinement, we successfully narrowed this list down to 37,960 high-quality, domain-specific concepts.
As an example, we show the extraction of concepts for the four papers used in Fig. 1:
1.
Accurate and rapid background estimation in single-molecule localization microscopy using the deep neural network BGnet [29]: ‘super resolution reconstruction’, ‘neural network’, ‘single molecule tracking’, ‘deep neural net’, ‘deep neural network’, ‘localization microscopy’, ‘biological structure’, ‘point source’, ‘point spread function’, ‘single molecule localization microscopy’, ‘optical microscopy’, and ‘neural net’.
2.
Machine learning for cluster analysis of localization microscopy data [30]: ‘neural network’, ‘supervised machine learning’, ‘spatial relation’, ‘localization microscopy’, ‘single molecule localization microscopy’, ‘neural net’, and ‘machine learning’.
3.
Constraints on cosmic strings using data from the first Advanced LIGO observing run [31]: ‘phase transition’, ‘cosmic string’, ‘gravitational wave’, ‘cosmic microwave’, ‘cosmic microwave background’, ‘topological defect’, ‘string theory’, and ‘ring theory’.
4.
Learning phase transitions from dynamics [32]: ‘neural network’, ‘recurrent network’, ‘time crystalline phase’, ‘phase transition’, ‘localization transition’, ‘spin chain’, ‘recurrent neural net’, ‘ct model’, ‘recurrent neural network’, ‘phase diagram’, ‘crystalline phase’, ‘neural net’, and ‘recurrent net’.
We created concept pairs, or edges, from the OpenAlex dataset, by detecting when domain-specific concepts co-occurred in paper titles or abstracts. This yielded 193,977,096 concept pairs across about 21 million papers. Each edge receives a time-stamp based on its paper’s publication date, converted to the number of days since January 1, 1990. The final full knowledge graph comprises 26,010,946 unique edges after merging multiple edges between the same concept pairs. The citation information for an edge includes the paper’s yearly citations from 2012 to 2023, alongside its total citation since publication. The OpenAlex dataset excludes yearly citations older than ten years, hence the focus on this specific ten-year time frame due to the absence of data prior to 2012. For edges formed by multiple papers, the edge weight combines the annual and total citations from all contributing papers.
Consider the edge formed by the concepts ‘single molecule localization microscopy’ and ‘neural net’, generated from paper [29] published on January 7, 2020. The time-stamp for this edge is derived from the days elapsed since January 1, 1990. The citation metrics for this edge includes the total and yearly citations from each contributing paper. Paper with 38 citations (=38), contributes yearly citations represented as for =2023,2022,…, 2012, with actual values for 2023 to 2020, and zeros for previous years, culminating in a citation sequence . Similarly, paper [30], published on March 20, 2020, with 43 citations (=43), adds its yearly citations for the same period (2023 to 2020). The aggregated citation data for this edge, combining and , yields a total of 81 citations, with an annual citation sequence of .
Training details – Our neural network consists of six fully connected layers, which include four hidden layers with 600 neurons each. The network inputs are 141 features for each pair of unconnected concepts , denoted as , where each . For instance, and represent the vertex degree of concepts and for the current year, . Detailed feature description and feature generation code are available in GitHub: Impact4Cast.
In our training process for the year 2016 to predict impact in 2019, we prepared a dataset comprising approximately 689 million unconnected concept pairs. The goal was to evaluate these pairs to determine whether their 3-year citation counts would have at least citations ( is impact range). From this extensive collection, we selected all positive samples (the 3-year citation counts are at least ). An equivalent number of negative samples were then randomly chosen to match the size of the positive set. The refined dataset was subsequently divided, allocating 85% for training and 15% for testing purposes. For the evaluation dataset in 2019, which aims to predict the impact by 2022, we randomly selected 10 million unconnected pairs. Our neural network was trained using the Adam optimizer with a learning rate of and a mini-batch size of 1000. In every training batch, we randomly chose an equal number of positive and negative samples from the training set. This approach was also applied to our 2019 training process for predictions into 2022, where the trained neural network is used for future forecasting.
The full dynamic knowledge graph, alongside the data required for feature preparation and evaluation, was processed using an Intel Xeon Gold 6130 CPU with 1 TiB RAM. The network training was performed on a NVIDIA Quadro RTX 6000 GPU, with 187 GiB RAM.
Individual feature’s predictive ability – In Fig. 5 (a), we observe an AUC score of 0.948 for the 2019 evaluation dataset with the neural network that uses all 141 features, trained on 2016 dataset and impact range . To explore the predictive ability of individual features, we trained separate neural networks on each feature using the same 2016 dataset, and then applied the 2019 evaluation dataset to these models. This resulted in 141 individual predictions, each from a network trained on a single feature. The features were ranked by their impact predictions, shown in the Extended Data Fig. 1.
Data availability
The data in this work are publicly available via Zenodo at https://doi.org/10.5281/zenodo.10692137 [55].
Code availability
The codes for this work can be found on the GitHub https://github.com/artificial-scientist-lab/Impact4Cast.
Acknowledgements
The authors thank the organizers of OpenAlex, arXiv, bioRxiv, and medRxiv, for making scientific resources freely and readily accessible. X.G acknowledges the support from the Alexander von Humboldt Foundation.